To precipitate all the sulfate ions from 22.0 mL of 0.150 molarity aluminum sulfate as barium sulfate, how many mL of 0.250 molarity3. barium nitrate are required? (8 points)




____Ba(NO3)2(aq)+____Al2(SO4)3(aq)→______BaSO4(s)+____Al(NO3)3(aq)













To neutralize 15.8 mL of 0.687 molarity sodium hydroxide, how much 0.376 molarity HCl is needed? (5 points)




____HCl(aq)+____NaOH(aq)→____NaCl(aq)+____H(OH)(l)






Glucose is a natural sugar that occurs in abundance. What is the mass percent of a solution formed from 148 g of glucose and 698 g of water? (5 points)












What is the mass percent sucrose in a solution made by combining 365 g of a 8.90 percent sucrose by mass aqueous solution with 145 g of an 6.80 percent sucrose by mass aqueous solution? (5 points)





1
Expert's answer
2022-03-27T10:30:46-0400

Moles of aluminium sulfate used=


(50/1000)× 0.350= 0.0175moles


Mole ratio= 3:1


Moles of barium nitrate=


0.0175× 3= 0.0525moles


(V/ 1000)× 0.250= 0.0525


V= 0.0525/0.00025 = 210ml